Shutdown tylan1 SNF 2008-05-01 13:02:06: repeated O2 flow faults

seymour at seymour at
Thu May 1 14:01:38 PDT 2008

checked and adjusted the O2HI mfc zero pot. Flowed O2 to setpoint of 2500 sccm and it was stable. I checked the remaining 3 mfc's and adjusted where necessary.

More information about the tylan1-pcs mailing list